# Structured Packing in LLM Training Improves Long Context Utilization

### v2

@@ Appendix B Detailed results
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline Repo
@@
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline Repo
@@
Table 9: To check the statistical significance of our results we prepare three subsets of C (details in Appendix F and train the models on a 50/50 mixture of RedPajama data (organized in standard way) and C data organized using one of the methods. Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
@@
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline Repo
@@
Table 10: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and other data organized using one of the methods. For details about training please refer to Section 3.1. Note that the model trained with SPLiCe on StackExchange outperforms the one trained on code on arXiv evaluation, showing the benefits of SPLiCe’s applicability to non-code data.
@@ C.1 SPLiCe parameters
There are two important design choices related to SPLiCe. First, how many related documents are retrieved in each step (the parameter $k$ in Algorithm 1). Second, how the documents are ordered. Table 12 indicates that $k=1$ is the best choice, though the differences are rather small. We found that changing the order of documents in training examples hardly matters. We use ’standard’, as ordered by Algorithm 1, the reversed order, and random shuffling. We consider exploring these design choices to be an interesting future work direction.
@@ C.2 Importance of separating tokens
We also evaluate the influence of BOS and EOS tokens on the performance. To be more precise, in all setups training examples are separated by BOS and EOS tokens. As both Repo and SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens. To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the dataset (sampling without replacement). The results are shown in Table 13. We note that the difference between the Baseline and the random concatenation approach is small and the random concatenation approach does not result in significant perplexity gains.
Table 13: Perplexity evaluation of two methods of organizing the data. Baseline – document equals training example. Random – concatenate documents into examples of length bounded by $120$k characters. Training examples are then fed into the model and separated by BOS and EOS tokens. The difference is negligible, which suggests that the extension of example length in Random does not help the model to utilize extended context. Experiments were performed using $270$M parameter model. We performed three runs on different subsets of C provide mean and standard deviation.
Training Method arXiv Code Code &
@@ Appendix D Key-Value retrieval task
Extract the value corresponding to the specified key in the JSON object below.
JSON data:
{"4ef217b7-6bc0-48c6-af35-2765f1e730f3": "068192b7-16b1-40e0-8495-61c63f979d50",
 "cd6b8bdc-bc6c-4490-acb4-bc187a2dccba": "7364a26e-289f-4968-93d3-b273e882bdee",
 "7d057372-4ab8-4811-8110-658c3f19fff4": "3ad075c5-b567-4201-85a7-cb31a0c91540",
 "c62e192d-45e6-4646-bb88-1529c73256c9": "f0411644-1f6d-42a6-8af8-f06da66efc77",
 "06134e93-e158-490e-a66c-8e3b98e12735": "50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",
 "3286f978-4270-4b54-8bfa-540d7e0772e6": "075cc716-1836-4f90-9be3-53e3d4ec6585",
 "4701aa05-c523-4b89-9700-64ab9c37c537": "49d86354-74c4-4256-9b3a-35e6e2b80d00",
 "c8895805-e574-4f13-9fe5-89da1d8c4748": "cc91af7f-8509-4bdc-bad7-2646af68e6d2"}
 "4701aa05-c523-4b89-9700-64ab9c37c537":
@@ Appendix E Perplexity Improvements
In Figure 5 we present perplexity improvements of SPLiCe BM25 over Baseline. Figure 6 shows evolution of SPLiCe model perplexity during the training. We follow (Anthropic, 2023) and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets. We average preplexity across arXiv, CUDA, Haskell and CommonCrawl datasets.
Figure 5: Perplexity improvement with SPLiCe against the Baseline of the final models (after $21$k training steps). We bucket tokens by their positions in the document, and calculate the average. Each dot is the difference of the averages of the SPLiCe and Baseline models. We observe that SPLiCe has smaller perplexity, and the improvements tend to be larger for tokens further in the document.
Figure 6: Evolution of the perplexity with SPLiCe, as the model is trained on more tokens. See 5 for the difference with the baseline. As expected, SPLiCe significantly improves perplexity for tokens whose positions are very distant in the sequence. Perplexity for more distant tokens improves more significantly compared to tokens in the beginning, early in the training.
Appendix F Data preparation
F.1 Evaluation Data
@@
When evaluating the model we skip documents that are shorter than $32$K tokens and truncate documents that are longer than that. Table 14 shows the number of tokens over which the perplexity was calculated.
Table 14: Number of evaluation tokens in each of the considered datasets.
Eval Dataset Number of Tokens
ArXiv 16M
C 16M
C# 16M
Common Lisp 16M
C++ 16M
CUDA 14M
Dart 16M
Emacs Lisp 15M
Erlang 16M
Fortran 16M
Go 16M
Groovy 4M
Haskell 16M
Java 16M
Pascal 16M
Python 16M
F.2 Train Data
The StackExchange data was taken from the Proof-pile. To prepare the code train data, we take the StarCoder train splits mentioned in Section F.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs. We have decided to divide the dataset in this way to not weaken the Repo method (a simple random split would scatter repository files across the data packs). We also split repos larger than $25$MB and filter out files that are longer than $30$k characters. The reason behind repo splitting is to avoid the situation where one repository occupies a significant portion of the data pack. We have noticed repos containing as many as $40$k files and files as long as $11$M characters. The character filtering is consistent with our method as we aim to improve the performance in a scenario that lacks high-quality long-context data. For C# and Python, only one pack is used for Baseline, Repo, and SPLiCe to organize the data. For C we have performed a run on three packs and provided results and standard deviation in Table 2. For the $3B$ model we run the methods on several packs and concatenate the results into a single dataset.
Appendix G TREC accuracy improvements
Figure 7 show details about accuracy improvements on TREC when considering different numbers of in-context examples.
(a) Examples: $190$, Context $4K$
(b) Examples: $380$, Context $8K$
(c) Examples: $780$, Context $16K$
@@ Appendix G TREC accuracy improvements
Figure 7: Histograms of accuracy improvement of SPLiCe BM25 over Baseline on TREC question classification task. The results are obtained by comparing the accuracy on the test set of TREC of the 3B model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples. Each set of in-context examples consists of elements randomly sampled (without replacement) from the training subset of TREC. Note that the model trained with SPLiCe is almost always better than the Baseline.
Appendix H Faiss parameters
Our experiments with SPLiCe Cont utilize Faiss (Johnson et al., 2017a) for fast approximate inner-product search. To be more precise we use the ”IVF8192,Flat” index that we train on 262144 examples coming from the dataset.

### v5

@@ Appendix B Detailed Results For Medium-Size Models
Eval/Method SPLiCeBM25 SPLiCeCont Baseline SPLiCeRepo
@@
Eval/Method SPLiCeBM25 SPLiCeCont Baseline SPLiCeRepo
@@
Table 11: To check the statistical significance of our results, we prepare three subsets of C (details in Appendix G and train the models on a 50/50 mixture of RedPajama data (organized in the standard way) and C data organized using one of the methods. Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
@@
Eval/Method SPLiCeBM25 SPLiCeCont Baseline SPLiCeRepo
@@
Table 12: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and other data organized using one of the methods. For details about training, please refer to Section 3.1. Note that the model trained with SPLiCe on StackExchange outperforms the one trained on code on arXiv evaluation, showing the benefits of SPLiCe’s applicability to non-code data.
@@
Table 13: Perplexity results comparing different ways of organizing the same data for non-FoT models. All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C# code is organized in one of three ways.
Altered Train Data: C#
Context $16$K: CodeLlama
Eval/Method SPLiCeBM25 Baseline SPLiCeRepo
ArXiv 5.74 5.76 5.74
C 2.66 2.70 2.66
C++ 2.79 2.83 2.79
CUDA 2.53 2.58 2.54
C# 1.91 1.93 1.91
Common Lisp 3.78 3.85 3.79
Dart 2.28 2.33 2.28
Emacs Lisp 8.29 8.41 8.30
Erlang 3.57 3.64 3.58
Fortran 3.93 4.01 3.95
Go 1.99 2.03 2.00
Groovy 2.95 3.01 2.96
Haskell 4.28 4.37 4.28
Java 2.31 2.35 2.31
Pascal 3.67 3.72 3.67
Python 3.22 3.27 3.22
Mean 3.49 3.55 3.50
Context $16$K: YaRN
Eval/Method SPLiCeBM25 Baseline SPLiCeRepo
ArXiv 5.77 5.79 5.77
C 2.68 2.72 2.68
C++ 2.81 2.85 2.81
CUDA 2.55 2.61 2.56
C# 1.92 1.94 1.92
Common Lisp 3.83 3.92 3.84
Dart 2.30 2.36 2.30
Emacs Lisp 8.37 8.52 8.38
Erlang 3.60 3.67 3.61
Fortran 3.97 4.06 3.99
Go 2.00 2.04 2.01
Groovy 2.98 3.03 2.98
Haskell 4.32 4.42 4.32
Java 2.33 2.37 2.33
Pascal 3.69 3.75 3.69
Python 3.24 3.29 3.24
Mean 3.52 3.58 3.53
Context $16$K: Naive
Eval/Method SPLiCeBM25 Baseline SPLiCeRepo
ArXiv 6.25 6.33 6.25
C 2.88 2.96 2.89
C++ 3.04 3.13 3.05
CUDA 2.84 2.96 2.85
C# 2.04 2.08 2.04
Common Lisp 4.40 4.56 4.39
Dart 2.50 2.60 2.51
Emacs Lisp 9.25 9.46 9.25
Erlang 3.98 4.10 4.00
Fortran 4.56 4.79 4.59
Go 2.14 2.21 2.16
Groovy 3.27 3.39 3.28
Haskell 4.84 5.03 4.87
Java 2.52 2.60 2.53
Pascal 4.05 4.20 4.10
Python 3.55 3.66 3.56
Mean 3.88 4.00 3.89
@@ C.1 SPLiCe Parameters
There are two important design choices related to SPLiCe. First, how many related documents are retrieved in each step (the parameter $k$ in Algorithm 1). Second, how the documents are ordered. Table 15 indicates that $k=1$ is the best choice, though the differences are rather small. We found that changing the order of documents in training examples hardly matters for $270$M models. We use ’standard’, as ordered by Algorithm 1, the reversed order, and random shuffling.
@@ C.2 Importance of Separating Tokens
We also evaluate the influence of BOS and EOS tokens on the performance. To be more precise, in all setups, training examples are separated by BOS and EOS tokens. As SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens. To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens, we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the dataset (sampling without replacement). The results are shown in Table 16. We note that the difference between the Baseline and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains.
Table 16: Perplexity evaluation of two methods of organizing the data. Baseline – document equals training example. Random – concatenate documents into examples of length bounded by $120$K characters. Training examples are then fed into the model and separated by BOS and EOS tokens. The difference is negligible, which suggests that the extension of example length in Random does not help the model to utilize extended context. Experiments were performed using the $270$M parameter model. We performed three runs on different subsets of C to provide mean and standard deviation.
Long Context Method arXiv Code Code &
@@ Appendix D Key-Value Retrieval Task
Extract the value corresponding to the specified key in the JSON object below. JSON data: {"4ef217b7-6bc0-48c6-af35-2765f1e730f3": "068192b7-16b1-40e0-8495-61c63f979d50",  "cd6b8bdc-bc6c-4490-acb4-bc187a2dccba": "7364a26e-289f-4968-93d3-b273e882bdee",  "7d057372-4ab8-4811-8110-658c3f19fff4": "3ad075c5-b567-4201-85a7-cb31a0c91540",  "c62e192d-45e6-4646-bb88-1529c73256c9": "f0411644-1f6d-42a6-8af8-f06da66efc77",  "06134e93-e158-490e-a66c-8e3b98e12735": "50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",  "3286f978-4270-4b54-8bfa-540d7e0772e6": "075cc716-1836-4f90-9be3-53e3d4ec6585",  "4701aa05-c523-4b89-9700-64ab9c37c537": "49d86354-74c4-4256-9b3a-35e6e2b80d00",  "c8895805-e574-4f13-9fe5-89da1d8c4748": "cc91af7f-8509-4bdc-bad7-2646af68e6d2"}  "4701aa05-c523-4b89-9700-64ab9c37c537":
We noted that FoT-trained models struggle with this task. This is probably due to the fact that they extend context only in a couple of layers, and the key-value retrieval task requires looking up and extracting a long sequence of letters and digits. Because of that, we evaluate FoT models with shorter dictionaries consisting of $75$ key-value pairs (around $6$K tokens) and show the results in Figures 3, and 3. For comparison, we also evaluate the $7$B CL model with this context length and show the results in Figure 3.
Figure 3: Performance on a smaller version of key-value retrieval task from (Liu et al., 2023). We note that FoT models (a), (b) generally struggle to retrieve tokens that are only visible to a subset of layers with extended context. For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama (Rozière et al., 2023) method of context extension. Each position was evaluated using $500$ examples.
@@ Appendix E Perplexity Improvements
In Figure 4 we present perplexity improvements of $3$B FoT SPLiCeBM25 over Baseline. Figure 5 shows the evolution of SPLiCe model perplexity during the training. We follow (Anthropic, 2023) and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets. We average perplexity across arXiv, CUDA, Haskell, and CommonCrawl datasets.
Figure 4: Perplexity improvement with SPLiCe against the Baseline of the final models (after $21$k training steps). We bucket tokens by their positions in the document and calculate the average. Each dot is the difference of the averages of the SPLiCe and Baseline models. We observe that SPLiCe has smaller perplexity, and the improvements tend to be larger for tokens further in the document.
Figure 5: Evolution of the perplexity with SPLiCe, as the model is trained on more tokens. See 4 for the difference with the baseline. As expected, SPLiCe significantly improves perplexity for tokens whose positions are very distant in the sequence. Perplexity for more distant tokens improves more significantly compared to tokens in the beginning, early in the training.
Appendix F Shuffling
Table 17: Average classification performance on TREC (Li & Roth, 2002; Hovy et al., 2001). We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1). SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was an identity function. We use $\pm$ to denote the standard deviation. We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
TREC
Model
Context length
(# of examples)
SPLiCe-no-shuf SPLiCe-shuf
$7$B CL $32$K (1560) 76.0 $\pm$2.5 76.6 $\pm$1.9
$8$K (380) 77.1 $\pm$2.4 76.5 $\pm$1.8
Appendix G Data Preparation
G.1 Evaluation Data
@@
When evaluating the perplexity of the model, we skip documents that are shorter than the model context and truncate documents that are longer than that. Table 18 shows the number of tokens over which the perplexity was calculated.
Table 18: Number of evaluation tokens in each of the considered datasets. For each context length $c$, we consider only documents having not less than $c$ tokens and extract the $c$ tokens prefix.
Eval/Method 16K 32K 64K
ArXiv 16M 16M 16M
C 16M 16M 16M
C++ 16M 16M 16M
CUDA 16M 14M 8M
C# 16M 16M 16M
Common Lisp 16M 16M 16M
Dart 16M 16M 16M
Emacs Lisp 16M 15M 9M
Erlang 16M 16M 12M
Fortran 16M 16M 16M
Go 16M 16M 16M
Groovy 11M 4M 2M
Haskell 16M 16M 15M
Java 16M 16M 16M
Pascal 16M 16M 16M
Python 16M 16M 16M
G.2 Train Data
The StackExchange data was taken from the Proof-pile. To prepare the code train data, we take the StarCoder train splits mentioned in Section G.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs. We also split repos larger than $25$MB and filter out files that are longer than $30$k characters. The reason behind repo splitting is to avoid the situation where one repository occupies a significant portion of the data pack. We have noticed repos containing as many as $40$K files and files as long as $11$M characters. The character filtering is consistent with our method as we aim to improve the performance in a scenario that lacks high-quality long-context data. For C# and Python, only one pack is used to organize the data. For C, we have performed a run on three packs and provided results and standard deviation in Table 3. For large models, we run the methods on several packs and concatenate the results into a single dataset. For natural language datasets, we extract a random subset of documents.
Appendix H Faiss Parameters
Our experiments with SPLiCeCont utilize Faiss (Johnson et al., 2019) for fast approximate inner-product search. To be more precise, we use the ”IVF8192,Flat” index that we train on 262144 examples coming from the dataset.
Appendix I Detailed Accuracy Improvements
The performance of in-context learning depends much on the choice of the in-context examples. To study this in more detail we study the following random variable
$\Delta(c)=\text{ACC}_{\textsc{{SPLiCe}}}(c)-\text{ACC}_{\textsc{Baseline}}(c),$
where $\text{ACC}_{\textsc{{SPLiCe}}}(c),\text{ACC}_{\textsc{Baseline}}(c)$ are the accuracies of the model trained with SPLiCe and Baseline respectively on the random choice of in-context examples $c$. Below we report the histograms of $\delta(c)$. In Table 4 and Table 5 we report mean $\Delta$ and $95\%$ confidence intervals as $\Delta$ [confidence interval].
Figure 6 shows details about accuracy improvements on TREC when considering different numbers of in-context examples.
@@ Appendix I Detailed Accuracy Improvements
Figure 6: Histograms of accuracy improvement of SPLiCeBM25 over Baseline on TREC question classification task. The results are obtained by comparing the accuracy on the test set of TREC of the $3$B FoT model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples. Each set of in-context examples consists of elements randomly sampled (without replacement) from the training subset of TREC. Note that the model trained with SPLiCe is almost always better than the Baseline.
(a) Examples: $190$, Context $16$K, Model: $3$B FoT
(b) Examples: $380$, Context $32$K, Model: $3$B FoT
(c) Examples: $190$, Context $16$K, Model: $7$B FoT
(d) Examples: $380$, Context $32$K, Model: $7$B FoT
(e) Examples: $190$, Context $16$K, Model: $7$B CL
(f) Examples: $380$, Context $32$K, Model: $7$B CL
Figure 7: Histograms of accuracy improvement of SPLiCeBM25 over Baseline on DBPedia. We sample $40$ sets of in-context examples and, for each set of in-context examples, evaluate on a random $500$ element subset of the DBPedia test set.
Appendix J Longer Context
In Table 19, we present the perplexity results for $270$M parameter models trained as described in Section 3 but with 2x larger context extension – context extended from $2$K to $64$K, instead of from $2$K to $32$K.
[…]
Long Context Method arXiv Code Code &
Data Haskell Python CUDA All arXiv
C# SPLiCeBM25 4.86 (.15) 2.60 (.19) 2.66 (.16) 2.32 (.19) 2.75 (.19) 2.88 (.19)
SPLiCeCont 4.88 (.13) 2.62 (.17) 2.67 (.15) 2.34 (.17) 2.77 (.17) 2.90 (.17)
SPLiCeRepo 4.88 (.13) 2.62 (.17) 2.68 (.14) 2.35 (.16) 2.77 (.17) 2.90 (.17)
Baseline 5.01 2.79 2.82 2.51 2.94 3.07
Appendix K Data Distribution Property
Chan et al. (2022); Han et al. (2023) shows that the ”burstiness” property of training data benefits a model’s generalisation. Specifically, ”burstiness” presents a flatter frequency distribution, with a relatively higher mass on the rare, long-tail tokens appearing in a sequence, which results in a lower Zipf’s coefficient of token frequency. We also observe a burstiness property of data in SPLiCe, as shown in Table 20.
Table 20: Zipf’s coefficient of token frequency on Baseline and SPLiCe. A lower Zipf’s coefficient represents a more significant burstiness property.
Training Data Context Length Method Zipf’s Coefficient
C 2K SPLiCeBM25 ${1.667}_{(0.139)}$
Baseline ${1.717}_{(0.152)}$
StackEx 2K SPLiCeBM25 ${1.836}_{(0.088)}$
Baseline ${1.865}_{(0.074)}$
C 32K SPLiCeBM25 ${1.512}_{(0.055)}$
Baseline ${1.593}_{(0.025)}$
StackEx 32K SPLiCeBM25 ${1.643}_{(0.026)}$
Baseline ${1.664}_{(0.013)}$